OpenStudy (anonymous):

Evaluate. Round to the nearest ten-thousandth. log5(4.7)

OpenStudy (anonymous):

log base 5 (4.7)

OpenStudy (anonymous):

Change of base formula.. \[\log_5{4.7} = \frac{\ln 4.7}{\ln 5}\]

OpenStudy (anonymous):

.96155465 what is the answer to the ten-thousandth

OpenStudy (anonymous):

4 places after the decimal...

OpenStudy (anonymous):

so would be .9616

OpenStudy (anonymous):

yes